# Integrating Text-to-Music Models with Language Models: Composing Long Structured Music Pieces

### v2

@@ Abstract
Recent music generation methods based on transformers have a context window of up to a minute. The music generated by these methods are largely unstructured beyond the context window. With a longer context window, learning long scale structures from musical data is a prohibitively challenging problem. This paper proposes integrating a text-to-music model with a large language model to generate music with form. We discuss our solutions to the challenges of such integration. The experimental results show that the proposed method can generate 2.5-minute-long music that is highly structured, strongly organized, and cohesive.
@@ 1 Introduction
Earlier works on generative music models, such as JukeBox, attempted to optimize the model to process long sequences of multiple minutes in the hope that the model would learn musical structures and forms at all scales. However, none of the models in the literature has demonstrated musical structure at a large scale and lack even simple musical forms. We argue that it is impractical to have a sufficiently large training dataset, as discussed in Section 2. Using Large Language Models (LLMs), we propose an approach to generate music with organized structures on larger scales. Specifically, we discuss our approach to integrate MusicGen with ChatGPT [11] in Section 4, after reviewing MusicGen briefly in Section 3. The experiments and evaluation of our method are presented in Section 5. Section 6 concludes the paper.
@@ 2 Learning Musical Form
Fig. 2: Illustrating the incoherence in images generated by Dall-E 344footnotemark: 4, Midjourney55footnotemark: 5, and Meta AI [^6] https://chatgpt.com/. These incoherences are evident in images featuring mirrors and wavering flags. Notice the forked or merged stripes on the flags and the inconsistent reflection and incidence angles in mirrors, among the other inconsistencies.
@@ 4 Controlling MusicGen by a Language Model
The first challenge is aligning the LLM with the text-to-music model. MusicGen has has been trained on the brief music description that are not technical and adhere to a certain style. The LLM needs to generate prompts that MusicGen can interpret. To bridge this gap, there are two main approaches.One is fine-tuning or training a pre-trained LLM, which involves adapting it directly to the task. Another approach is in-context learning, which offers the advantages of requiring fewer samples and being less resource-intensive when trying and evaluating different prompts. Notably, using a large number of samples for in-context learning has been shown to outperform fine-tuning in terms of accuracy [21, 22].
We use in-context learning to instruct ChatGPT to generate prompts for MusicGen, with providing 50 song descriptions from Pond5. To find the required number of samples we observer how often the generated music is faithful to the text prompt to MusicGen. Our empirical results show with about 10 song descriptions the generated by prompts were often misinterpreted by MusicGen. Increasing the number of samples from 50 to 80 showed no improvement in the interpretability of the prompts by MusicGen. As LLMs are more prone to hallucination [23], more than necessary samples should be avoided.
ChatGPT tends to mix multiple genres in the prompts for a single piece, resulting in structures that are not particularly well-designed. However, when we direct ChatGPT to consider a framework such as the ITPRA theory [24], the resulting structures are more organized and coherent. Additionally, employing the chain of thought approach [25]—asking it to first respond with a description of a song and its form, followed by generating prompts for the parts—further improves the organization of the musical structure. We have also implemented additional rules in the prompts to ChatGPT to introduce constraints specific to MusicGen. For instance, since MusicGen cannot generate parts that have ”a slower tempo” than a previous part, we instruct ChatGPT to generate the prompts, the length of each part in seconds, and the musical elements from prior parts that should be used, all formatted in JSON. In the following prompt excerpt, some sections have been omitted for brevity.
@@ 5 Evaluation
Switching between prompts passed to MusicGen creates a sudden jump. To ensure a smooth transition between parts, we modify the CFG method. Instead of estimating one conditioned probability distribution, we estimate two conditioned distributions. In the interpolated probability distribution, the weight of the first conditioned distribution over five seconds goes from one to zero linearly, and the weight of the second conditioned distribution increases from zero to one. This method allows the model to generate tokens that facilitate a continuous transition between parts.
@@
Similarly, to generate parts with variations of the prior parts, we provide a 15-second audio prompt from a previous part to MusicGen. This 15-second audio is passed to the encoder of EnCodec to generate the corresponding tokens. Then, the prompt tokens are passed to the predictive model of MusicGen. The weight of the probability distribution over tokens, conditioned on these audio prompt tokens, is then linearly reduced to zero over 10 seconds.
@@
Figure 3 compares the self similarity matrices (SSM) of samples from MusicGen, our method, and Pond5. We use the fusion method proposed in  [26] to generate the combined SSMs. The SSMs of the sample from our method resemble the variation and similarity structure in the sample from Pond5. The
@@
After downsampling the fused SS matrices of 100 samples, the mean and variance matrices are estimated for Pond5, our method, and MusicGen in Figure 1. Therefore, this figure helps to compare the distributions of the music structures. It is apparent that, from the mean and variation values from the top-right of the matrices, the MusicGen samples do not have parts near the end of each piece that resemble the parts near the beginning. In contrast, the samples from our method are more similar to the samples from Pond5. We estimate the Fréchet distance between the two pairs of distributions, but instead of using Inception feature vectors [27], we estimate the mean and covariance matrices of the upper triangle of the fused SS matrices. The Fréchet distances between the the distributions of the Pond5 samples and ours is 0.086, and between the Pond5 and MusicGen samples is 0.108, supporting the claim the structure of the samples generated by our method are more similar to the samples composed by musicians from Pond5.
@@
For a subjective evaluation by non-musicians, we generate 10 samples using our method, 10 samples using MusicGen, and 10 samples from Pond5. Each of these samples is 2.5 minutes long. Using Amazon Mechanical Turk (MTurk), we collect a mean opinion score (MOS) between 1 and 5 for each sample from 10 non-musicians. We ask the human raters to evaluate the overall quality of the music, following the recommended practices in CrowdMOS [28]. The subjects are told a score of 1 means they dislike the music and would not like to listen to similar music. A score of 5 means they find the music interesting and would like to listen to similar music. Given long music tracks with more organized structure are expected to be more engaging, we use the likeness of the samples as a proxy for the improved structure and form. The results in Figure 4(top) indicate that adding musical form through our method to MusicGen improves the perceived quality of the music, almost on par with human-composed music pieces from Pond5.
@@ 6 Conclusions
This paper argues that due to the discussed nature of the problem, the music generative models cannot learn long-scale structure or musical form from a musical datasets. We then present a novel method to combine music generative models with LLMs to generate music with a musical form. The technical challenges are discussed. The subjective and objective evaluations support the claim that our method can generate well-structured 2.5-minute music.

### v3

@@ Abstract
Recent music generation methods based on transformers have a context window of up to a minute. The music generated by these methods is largely unstructured beyond the context window. With a longer context window, learning long-scale structures from musical data is a prohibitively challenging problem. This paper proposes integrating a text-to-music model with a large language model to generate music with form. The papers discusses the solutions to the challenges of such integration. The experimental results show that the proposed method can generate 2.5-minute-long music that is highly structured, strongly organized, and cohesive.
@@ 1 Introduction
Earlier works on generative music models, such as JukeBox, attempted to optimize the model to process long sequences of multiple minutes in the hope that the model would learn musical structures and forms at all scales. However, none of the models in the literature has demonstrated musical structure at a large scale and lack even simple musical forms. Section 2 presents the arguement that it is impractical to have a sufficiently large training dataset. Using Large Language Models (LLMs), this papers proposes an approach to generate music with organized structures on larger scales. Section 3 presents a brief review of MusicGen. Sections 4 discusses the proposed approach to integrate MusicGen with ChatGPT [11]. The experiments and evaluation of our method are presented in Section 5. Section 6 concludes the paper.
@@ 2 Learning Musical Form
Fig. 2: Illustrating the incoherence in images generated by Dall-E 344footnotemark: 4, Midjourney55footnotemark: 5, and Meta AI [^6] https://chatgpt.com/. These inconsistencies are evident in images featuring mirrors and wavering flags. Notice the forked or merged stripes on the flags and the inconsistent reflection and incidence angles in mirrors, among the other inconsistencies.
@@ 4 Controlling MusicGen by a Language Model
The first challenge is aligning the LLM with the text-to-music model. MusicGen has has been trained on brief music descriptions that are not technical and adhere to a certain style. The LLM needs to generate prompts that MusicGen can interpret. To bridge this gap, there are two main approaches.One is fine-tuning or training a pre-trained LLM, which involves adapting it directly to the task. Another approach is in-context learning, which offers the advantages of requiring fewer samples and being less resource-intensive when trying and evaluating different prompts. Notably, using a large number of samples for in-context learning has been shown to outperform fine-tuning in terms of accuracy [21, 22].
In-context learning is used to instruct ChatGPT to generate prompts for MusicGen by providing 50 song descriptions from Pond5. To find the required number of samples, it is estimated how often the generated music is faithful to the text prompt to MusicGen. The empirical results show that with about 10 song descriptions the generated by prompts were often misinterpreted by MusicGen. Increasing the number of samples from 50 to 80 showed no improvement in the interpretability of the prompts by MusicGen. As LLMs are more prone to hallucination [23], more than necessary samples should be avoided.
ChatGPT tends to mix multiple genres in the prompts for a single piece, resulting in structures that are not particularly well-designed. However, when ChatGPT is directed to consider a framework such as the ITPRA theory [24], the resulting structures are more organized and coherent. Additionally, employing the chain of thought approach [25]—asking it to first respond with a description of a song and its form, followed by generating prompts for the parts—further improves the organization of the musical structure. Additional rules are added in the prompt to ChatGPT to introduce constraints specific to MusicGen. For instance, since MusicGen cannot generate parts that have ”a slower tempo” than a previous part, ChatGPT is instructed to generate the prompts, the length of each part in seconds, and the musical elements from prior parts that should be used, all formatted in JSON. In the following prompt excerpt, some sections have been omitted for brevity.
@@ 5 Evaluation
Switching between prompts passed to MusicGen creates a sudden jump. To ensure a smooth transition between parts, the CFG method is modified. Instead of estimating one conditioned probability distribution, two conditioned distributions are estimated. In the interpolated probability distribution, the weight of the first conditioned distribution over five seconds goes from one to zero linearly, and the weight of the second conditioned distribution increases from zero to one. This method allows the model to generate tokens that facilitate a continuous transition between parts.
@@
Similarly, to generate parts with variations of the prior parts, a 15-second audio prompt from a previous part is provided to MusicGen. This 15-second audio is passed to the encoder of EnCodec to generate the corresponding tokens. Then, the prompt tokens are passed to the predictive model of MusicGen. The weight of the probability distribution over tokens, conditioned on these audio prompt tokens, is then linearly reduced to zero over 10 seconds.
@@
Figure 3 compares the self similarity matrices (SSM) of samples from MusicGen, our method, and Pond5. The fusion method proposed in  [26] is used to generate the combined SSMs. The SSMs of the sample from our method resemble the variation and similarity structure in the sample from Pond5.
@@
After downsampling the fused SS matrices of 100 samples, the mean and variance matrices are estimated for Pond5, our method, and MusicGen in Figure 1. Therefore, this figure helps to compare the distributions of the music structures. It is apparent that, from the mean and variation values from the top-right of the matrices, the MusicGen samples do not have parts near the end of each piece that resemble the parts near the beginning. In contrast, the samples from our method are more similar to the samples from Pond5. The Fréchet distance between the two pairs of distributions is estimated, but instead of using Inception feature vectors [27], the mean and covariance matrices of the upper triangle of the fused SS matrices are used in the distance estimation. The Fréchet distances between the the distributions of the Pond5 samples and ours is 0.086, and between the Pond5 and MusicGen samples is 0.108, supporting the claim the structure of the samples generated by our method are more similar to the samples composed by musicians from Pond5.
@@
For a subjective evaluation by non-musicians, 10 samples are generated using our method, 10 samples using MusicGen, and 10 samples from Pond5. Each of these samples is 2.5 minutes long. Using Amazon Mechanical Turk (MTurk), a mean opinion score (MOS) between 1 and 5 for each sample from 10 non-musicians is collected. The human raters are asked to evaluate the overall quality of the music, following the recommended practices in CrowdMOS [28]. The subjects are told a score of 1 means they dislike the music and would not like to listen to similar music. A score of 5 means they find the music interesting and would like to listen to similar music. Given long music tracks with more organized structure are expected to be more engaging, the likeness of the samples is used as a proxy for the improved structure and form. The results in Figure 4(top) indicate that adding musical form through our method to MusicGen improves the perceived quality of the music, almost on par with human-composed music pieces from Pond5.
@@ 6 Conclusions
This paper argues that due to the discussed nature of the problem, the music generative models cannot learn long-scale structure or musical form from a musical datasets. Then a novel method to combine music generative models with LLMs to generate music with a musical form is presented. The technical challenges are discussed. The subjective and objective evaluations support the claim that our method can generate well-structured 2.5-minute music.